The temperature of an oven is kept constant at 1050 K. A hole with a diameter of 21.0 mm is drilled in the wall of the oven. How much power is emitted by this hole? (Consider the hole as a black body.) What is the wavelength for which the radiant energy is maximum?
Answer #1


T = 1050 K, d = 21 mm

P = (sigma)*AT^4

sigma =5.67*10^-8 W/m^2.K^4

P = (5.67*10^-8)(3.14*(0.021/2)^2)*(1050)^4

P = 23.86 W

From weins displacment law

(lamda)*T = 2.89*10^-3 m.K

(lamda)*1050 = 2.89*10^-3

lamda = 2.75 um